# The Christoffel problem for the disk area measure

### v2

@@ Abstract.
The mixed Christoffel problem asks for necessary and sufficient conditions for a Borel measure on the Euclidean unit sphere to be the mixed area measure of some convex bodies, all but one of them are fixed. We provide a solution in the case where the fixed reference bodies are (n1)(n-1)-dimensional disks with parallel axes and the measure has no mass at the poles of the sphere determined by this axis.
@@ 1. Introduction
For a convex body (convex, compact set) $K\subset\mathbb{R}^{n}$ with smooth boundary, the area measure of order $1$, denoted $S_{1}(K,\cdot)$, is the absolutely continuous measure on the unit sphere $\mathbb{S}^{n-1}$ with density given by the mean radius of curvature of $K$ as a function of the outer unit normal vector. Extending this definition, the definition of $S_{1}(K,\cdot)$ can be extended to general convex bodies via variations of the surface area measure $S_{n-1}(K,\cdot)$, that is,
@@
Dating back to the nineteenth century with the work of Christoffel [Christoffel1865], it was an important question in convex geometry, nowadays known as the Christoffel problem, to determine the class of Borel measures on the sphere that arise as $S_{1}(K,\cdot)$ for some convex body $K$. After Christoffel’s work [Christoffel1865], treating the three dimensional case, a complete solution was obtained at the end of the 1960s with the works of Berg [Berg1969] and Firey [Firey1967, Firey1968]. Special cases were treated for smooth bodies by Pogorelov [Pogorelov1953] and for polytopes by Schneider [Schneider1977b].
The key to the solution of the Christoffel problem was the observation that $S_{1}(K,\cdot)$ can be obtained from the support function $h_{K}$ of $K$ by the means of the differential operator
$\Box_{n}=\frac{1}{n-1}\Delta_{\mathbb{S}}+\mathrm{Id},$
where $\Delta_{\mathbb{S}}$ is the spherical laplacian on $\mathbb{S}^{n-1}$. Indeed, $S_{1}(K,\cdot)=\Box_{n}h_{K}$ in the distributional sense, and using spherical harmonics, Berg [Berg1969] constructed a family of functions $g_{n}$, $n\geq 2$, on (1,1)(-1,1) that serve as integral kernels of the inverse operator to $\Box_{n}$. Writing $\mathcal{K}(\mathbb{R}^{n})$ for the set of convex bodies in $\mathbb{R}^{n}$, Berg’s result [Berg1969]*Thm. 5.3 can be formulated as follows.
@@
(1.2) $h(u)=\int_{\mathbb{S}^{n-1}}g_{n}(\langle u,v\rangle)\,\mu(dv),\qquad u\in\mathbb{S}^{n-1},$
@@
Here and in the following, a measure $\mu$ on $\mathbb{S}^{n-1}$ is called centered if $\int_{\mathbb{S}^{n-1}}ud\mu(u)=0$. Moreover, we denote by $\langle\cdot,\cdot\rangle$ the euclidean inner product and by
(1.3) $s(K)=\frac{1}{\kappa_{n}}\int_{\mathbb{S}^{n-1}}h_{K}(u)u\,du\in\mathbb{R}^{n}$
the Steiner point of a given convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ (see, e.g., [Schneider2014]*p. 50).
Let us spell out the explicit formula for the function $g_{2}$ for later reference:
$g_{2}(t)=\sqrt{1-t^{2}}(\pi-\arccos t)+ct,$
with $c$ chosen in such a way to make the function $u\mapsto g_{2}(\langle e_{n},u\rangle)$ on $\mathbb{S}^{n-1}$ centered.
In this note, we consider an anisotropic version of $S_{1}(K,\cdot)$, obtained by replacing the unit ball $B^{n}$ in ˜1.1 by an arbitrary convex body $C\in\mathcal{K}(\mathbb{R}^{n})$
(1.4) $S_{1}(K,C;\beta)=\frac{1}{(n-1)!}\left.\left(\frac{d}{dt}\right)^{\!n-2}\right|_{t=0^{+}}S_{n-1}(K+tC,\beta),$
for every Borel set $\beta\subseteq\mathbb{S}^{n-1}$, which gives rise to the anisotropic Christoffel problem:
@@
This problem has only started to gain attention very recently (e.g., [Colesanti2025a]), and very little is known about it. Let us note that it was also considered in the broader range of an anisotropic Christoffel–Minkowski problem, that is, for intermediate area measures, in [Brauner2025a], where it was solved in the case where all bodies are bodies of revolution. Indeed, this note originated from a side observation during the project [Brauner2025a].
1.1. Main result
Our main result is a solution of the anisotropic Christoffel problem in the case where $C=\mathbb{D}$, that is, for the disk area measure $S_{1}(K,\mathbb{D};{{}\cdot{}})$, under an additional condition on the mass at the north and south pole $\pm e_{n}\in\mathbb{S}^{n-1}$. We hope that, in analogy to the proof strategy in [Brauner2025a], this can serve as an initial step to solve the general anisotropic Christoffel problem via transforming the area measures.
In order to state our theorem, let $\pi:\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}\to\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$ denote the map that assigns to every $u\in\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$ the unique linear $2$-space $e_{n}\vee u$ containing both $u$ and the $n$-th coordinate vector $e_{n}$. Here, we denote by $\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$ the grassmanian of all $2$-dimensional subspaces of $\mathbb{R}^{n}$ that contain $e_{n}$.
Under the assumption that the pushforward $\pi_{\ast}\mu$ is absolutely continuous with continuous density, the measure $\mu$ disintegrates into a family $(\mu_{E})_{E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}$ of measures on $S^{1}(E)$, such that
@@ 1.1. Main result
Note that, using the condition that $\mu(\{\pm e_{n}\})=0$, we tacitly identify $\mu$ with its restriction to $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$. See Section˜2.2 for the details on the construction of this disintegration.
@@
    (1.6) $h(u)-\frac{1}{\pi}\int_{\mathbb{S}^{1}(e_{n}\vee u)}h(v)\langle u,v\rangle\,dv=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}(\pi-\arccos\langle u,v\rangle)\,\mu_{e_{n}\vee u}(dv),$
@@
Let us comment on the condition that the measure has no mass at the poles $\pm e_{n}$. This is due to the fact that the solution of the Christoffel problem for the disk is not unique as soon as there is a face in either direction $\pm e_{n}$. Indeed, for every $K\in\mathcal{K}(e_{n}^{\perp})$, $S_{1}(K,\mathbb{D},\cdot)=V^{e_{n}^{\perp}}(K,\mathbb{D}^{[n-2]})(\delta_{e_{n}}+\delta_{-e_{n}})$ only depends on the mean width of $K$.
On a technical level, mass at the poles $\pm e_{n}$ raises the problem of how to define a disintegration similar to ˜1.5. Indeed, the domains of the measures $\mu_{E}$ would overlap in the poles $\pm e_{n}$, whence a method is needed to distribute the mass $\mu(\{\pm e_{n}\})$ among the measures $\mu_{E}$. However, this corresponds to determine the (highly non-unique) shape of the faces of $K$ in direction $\pm e_{n}$. As soon as a disintegration satisfying ˜1.6 is constructed, the steps in the proof of Theorem˜A can be repeated to yield the solution.
2. Proof of the main result
Our proof is inspired by the the Kubota-type formula of Hug, Mussnig, and Ulivelli [Hug2024] in the instance where $i=1$: For a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ and a bounded Borel function $f:\mathbb{S}^{n-1}\to\mathbb{R}$,
@@ 2. Proof of the main result
This integral geometric formula expresses the mixed area measure $S_{1}(K,\mathbb{D},{}\cdot{})$ as an average of surface area measures of two-dimensional projections of $K$. The core idea of our argument is to compare ˜2.1 to the disintegration ˜1.5 of the measure $\mu$ to relate the anisotropic Christoffel problem with the disk to the classical Christoffel problem in two dimensions.
By ˜1.1, we explicitly solve the $2$-dimensional Christoffel problem for the measures $\mu_{E}$, giving bodies $K_{E}\in\mathcal{K}(E)$, $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. If there is a body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that every $K_{E}$ arises as the orthogonal projection $K|E$ of $K$ onto $E$, then we can conclude $\mu=S_{1}(K,\mathbb{D};{{}\cdot{}})$. However, as the subspaces $E$ intersect only in the line spanned by $e_{n}$, compatibility of the $K_{E}$ is no issue; convexity is asserted by condition ˜(iii) in Theorem˜A.
@@
In the following, we first recall some classical facts from convex geometry, in particular on mixed volumes and mixed area measures that will be needed later on. Then we show the existence of the disintegration ˜1.5. In the last section we finally prove Theorem˜A.
@@ 2.1. Preliminaries
First, recall that the mixed volume $V(K_{1},\dots,K_{n})$, $K_{1},\dots,K_{n}\in\mathcal{K}(\mathbb{R}^{n})$ is defined as the suitable coefficient of the homogeneous polynomial obtained by Minkowski addition
@@
with $F(K,u)=\{x\in\mathbb{R}^{n}:h_{K}(u)=\langle x,u\rangle\}$ the face of $K$ in direction $u\in\mathbb{S}^{n-1}$.
@@
In order to deal with the uniqueness of the solution $K$ to the Christoffel problem involving the disk, we recall Minkowski’s quadratic inequality, a classical result from Brunn–Minkowski theory. It states that for all convex bodies $K,L,C\in\mathcal{K}(\mathbb{R}^{n})$,
@@
Let $K,L,C\in\mathcal{K}(\mathbb{R}^{n})$ be such that $C-C\subseteq e_{n}^{\perp}$ and $V(K,L,C^{[n-2]})>0$. Then equality holds in ˜2.5 if and only if $\tilde{L}:=\frac{V(K,L,C^{[n-2]})}{V(L,L,C^{[n-2]})}L$ satisfies that $K+F(\tilde{L},e_{n})$ and $\tilde{L}+F(K,e_{n})$ have the same supporting hyperplanes in all $1$-extremal normal directions of $C$.
Observe now that at every point $x\in\mathbb{D}$, the corresponding normal cone $N(\mathbb{D},x)$ is either one- or two-dimensional. This implies that every direction $u\in\mathbb{S}^{n-1}$ is a $1$-extremal normal direction of $\mathbb{D}$, so ˜2.2 specializes for $C=\mathbb{D}$ to the following statement.

### v3

@@ Abstract.
The mixed Christoffel problem asks for necessary and sufficient conditions for a Borel measure on the Euclidean unit sphere to be the mixed area measure of some convex bodies, all but one of them are fixed. We consider the case in which the reference bodies are (n1)(n-1)-dimensional disks lying in a fixed hyperplane. We obtain an integral representation that reconstructs the support function of a convex body from its disk area measure, without any regularity assumptions. In the smooth setting, we reformulate the problem as a linear differential equation on the sphere, and derive a necessary and sufficient condition on the density of the disk area measure guaranteeing both convexity and regularity of the solution.
@@ 1. Introduction
For a convex body (convex, compact set) $K\subset\mathbb{R}^{n}$ with smooth boundary, the first-order area measure, denoted $S_{1}(K,\cdot)$, is the absolutely continuous measure on the unit sphere $\mathbb{S}^{n-1}$ with density given by the mean radius of curvature of $K$ as a function of the outer unit normal vector. The definition of $S_{1}(K,\cdot)$ can be extended to general convex bodies via variations of the surface area measure $S_{n-1}(K,\cdot)$, that is,
@@
Dating back to the nineteenth century with the work of Christoffel [Christoffel1865], it was an important question in convex geometry, nowadays known as the Christoffel problem, to determine the class of Borel measures on the sphere that arise as $S_{1}(K,\cdot)$ for some convex body $K$. After Christoffel’s work [Christoffel1865], treating the three dimensional case, a complete solution was obtained at the end of the 1960s with the works of Berg [Berg1969] and Firey [Firey1967, Firey1968] (see also [Li2021] for a more recent approach). Special cases were treated for smooth bodies by Pogorelov [Pogorelov1953] and for polytopes by Schneider [Schneider1977b].
The key to solving the Christoffel problem was the observation that in the sense of distributions on $\mathbb{S}^{n-1}$, the area measure $S_{1}(K,\cdot)$ can be obtained from the support function $h_{K}$ of $K$ by means of a differential operator:
(1.2) $S_{1}(K,{}\cdot{})=\frac{1}{n-1}\Delta_{\mathbb{S}}h_{K}+h_{K},$
where $\Delta_{\mathbb{S}}$ denotes the Laplace–Beltrami operator on $\mathbb{S}^{n-1}$. Using spherical harmonics, Berg [Berg1969] constructed a family of functions $g_{n}\in C^{\infty}(-1,1)$, $n\geq 2$, that serve as Green functions of the operator. Here and in the following, we denote by $\mathcal{K}(\mathbb{R}^{n})$ the set of convex bodies in $\mathbb{R}^{n}$, by $s(K)$ the Steiner point of a body $K\in\mathcal{K}(\mathbb{R}^{n})$ (see Section˜2.1), and call a Borel measure $\mu$ on $\mathbb{S}^{n-1}$ centered if $\int_{\mathbb{S}^{n-1}}u\,\mu(du)=o$. Berg’s result [Berg1969]*Thm. 5.3 can be stated as follows.
@@
(1.3) $h(u)=\int_{\mathbb{S}^{n-1}}g_{n}(\left\langle u,v\right\rangle)\,\mu(dv),\qquad u\in\mathbb{S}^{n-1},$
@@
Recently, more general versions of the Christoffel problem, involving mixed area measures, have attracted considerable attention. Mixed area measures are closely related to mixed volumes of convex bodies, both of which are fundamental notions in convex geometry (see, e.g. [Schneider2014]). The first-order area measure $S_{1}(K,\cdot)$ of a body $K\in\mathcal{K}(\mathbb{R}^{n})$ is an instance of this. Replacing the unit ball $B^{n}$ in ˜1.1 by an arbitrary convex body $C\in\mathcal{K}(\mathbb{R}^{n})$, we obtain a rich family of mixed area measures:
(1.4) $S_{1}(K,C;\beta)=\frac{1}{(n-1)!}\left.\left(\frac{d}{dt}\right)^{\!n-2}\right|_{t=0^{+}}S_{n-1}(K+tC,\beta)$
for every Borel set $\beta\subseteq\mathbb{S}^{n-1}$. In this article, we consider the following Christoffel type problem.
@@
This is an instance of the mixed Christoffel problem, where the role of the reference body $C$ is taken by a collection of reference bodies $C_{1},\ldots,C_{n-2}$. For bodies that are symmetric about an axis, this problem has been completely solved by the authors of this article [Brauner2025a], in the spirit of the classical work of Firey [Firey1970]. The corresponding problem for radially symmetric convex functions and mixed Monge–Ampère measures was treated by Mussnig–Ulivelli [Mussnig2025]. The two approaches are independent of each other, although both heavily exploit the imposed symmetry.
In the general setting, recently, a significant advancement was achieved by Colesanti–Focardi–Guan–Salani [Colesanti2025], who identified sufficient conditions characterizing a rich class of mixed area measures. Their approach is based on the techniques of Guan–Ma [Guan2003] and relies on smoothness assumptions. To date, the only instance of the mixed Christoffel problem where a complete solution is known, including necessary and sufficient conditions, remains the classical Christoffel problem (where $C=B^{n}$). While the solution of Berg and Firey is somewhat formal, the work of Li–Wan–Wang [Li2021] provides a more readily checkable condition via the fundamental solution to the Laplace equation.
Main results
The aim of this article is to fully resolve another instance of the mixed Christoffel problem, namely where $C$ is the (n1)(n-1)-dimensional unit disk perpendicular to the $n$-th coordinate vector $e_{n}$, denoted as $\mathbb{D}$. Such disk area measures involving the disk $\mathbb{D}$ turn out to possess particularly nice geometric properties, and have proved useful in a variety of applications (see [Brauner2024a, Brauner2025a, Hug2024, Hug2024a, Mussnig2025, Colesanti2022a]). In particular, for axially symmetric bodies, the authors of this article reduce the mixed Christoffel problem (and more generally, the mixed Christoffel–Minkowski problem) to the one involving the disk area measure [Brauner2025a].
Our contribution to this problem is twofold. First, in the same spirit as Berg and Firey, we establish an integral representation expressing the support function of $K$ in terms of the disk area measure $S_{1}(K,\mathbb{D};\cdot)$, without any regularity assumptions on $K$. Second, we treat the problem as a differential equation and derive an efficient condition on the density of the disk area measure to determine both the convexity and regularity of the solution.
In order to state our first main result, we let $\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$ denote the Grassmanian of all two-dimensional linear subspaces of $\mathbb{R}^{n}$ containing the $n$-th coordinate vector $e_{n}$, and let $\pi:\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}\to\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$ be the map that assigns to every $u\in\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$ the unique plane $e_{n}\vee u$ spanned by $u$ and $e_{n}$. Given a Borel measure $\mu$ on $\mathbb{S}^{n-1}$, if its pushforward $\pi_{\ast}\mu$ is absolutely continuous with continuous density, then the measure $\mu$ disintegrates into a family $(\mu_{E})_{E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}$ of measures on $\mathbb{S}^{1}(E)$ such that for any function $f\in C(\mathbb{S}^{n-1})$,
@@
Note that, as long as $\mu(\{\pm e_{n}\})=0$, we may identify $\mu$ with its restriction to $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$; see Section˜2.2 for the details on the construction of this disintegration.
@@
    (1.6) $h(u)-\frac{1}{\pi}\int_{\mathbb{S}^{1}(e_{n}\vee u)}h(v)\left\langle u,v\right\rangle\,dv=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\left\langle u,v\right\rangle^{2}}(\pi-\arccos\left\langle u,v\right\rangle)\,\mu_{e_{n}\vee u}(dv),$
@@
Let us comment on the condition that the measure has no mass at the poles $\pm e_{n}$. This is due to the fact that the solution of the Christoffel problem for the disk is not unique as soon as there is a face in either direction $\pm e_{n}$. Indeed, whenever $K\in\mathcal{K}(e_{n}^{\perp})$, its disk area measure $S_{1}(K,\mathbb{D},\cdot)=V^{e_{n}^{\perp}}(K,\mathbb{D}^{[n-2]})(\delta_{e_{n}}+\delta_{-e_{n}})$ only depends on the mean width of $K$.
On a technical level, mass at the poles $\pm e_{n}$ raises the problem of how to define a disintegration similar to ˜1.5. Indeed, the domains of the measures $\mu_{E}$ overlap at the poles $\pm e_{n}$, so a procedure is required to distribute the mass $\mu(\{\pm e_{n}\})$ among the measures $\mu_{E}$. However, this amounts to determining the (highly non-unique) geometry of the faces of $K$ in the directions $\pm e_{n}$. As soon as a disintegration satisfying ˜1.6 is constructed, the steps in the proof of ˜A can be repeated to yield the solution.
Let us emphasize that ˜A provides a set of conditions that are necessary and sufficient without requiring any smoothness of $K$, and that it allows to retrieve the support function of $K$ from its disk area measure. However, as in the work of Berg and Firey, it does not offer a practical way to check convexity or to determine the regularity of $K$ depending on the regularity of $\mu$. Hence we also pursue a second approach, where we consider the problem as a linear differential equation, similar to ˜1.2.
If we assume $K\in\mathcal{K}(\mathbb{R}^{n})$ to be of class $C^{2}_{+}$ (that is, its boundary is a $C^{2}$ submanifold of $\mathbb{R}^{n}$ with everywhere positive Gauss curvature), then its support function $h_{K}$ is a $C^{2}(\mathbb{S}^{n-1})$ function. Parametrizing $\mathbb{S}^{n-1}$ in polar coordinates $u=\cos\theta e_{n}+\sin\theta w$, where $\theta\in[0,\pi]$ and $w\in\mathbb{S}^{n-2}=\mathbb{S}^{n-2}(e_{n}^{\perp})$, the disk area measure $S_{1}(K,\mathbb{D},{}\cdot{})$ can be expressed in terms of $h_{K}$ as
(1.7) $S_{1}(K,\mathbb{D},du)=\frac{1}{n-1}\big(\partial_{\theta}^{2}h_{K}(\theta,w)+h_{K}(\theta,w)\big)d\theta\,dw.$
Notably, the lower dimensionality of the disk $\mathbb{D}$ causes the associated partial differential equation to degenerate into a family of ordinary differential equations along the meridians of the unit sphere, allowing us to employ standard theory of ordinary differential equations.
As in ˜A, the core of the argument consists in gluing together the solutions constructed on the meridians to obtain a global convex solution on $\mathbb{S}^{n-1}$. The advantage of working in the class $C^{2}_{+}$ is that the convexity of the resulting solution $K$ can be encoded as the positive definiteness of the Hessian of its support function. Our second main result provides a practical condition that fully characterizes disk area measures of $C^{2}_{+}$ regular convex bodies.
Theorem B.
Let $q\in C([0,\pi]\times\mathbb{S}^{n-2})$ be strictly positive and $q(\theta,\,\cdot\,)\in C^{2}(\mathbb{S}^{n-2})$ for all $\theta\in[0,\pi]$. There exists some convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ of class $C^{2}_{+}$ such that $q(\theta,w)d\theta\,dw=S_{1}(K,\mathbb{D},du)$ if and only if there exist $t\in\mathbb{R}$ and $x\in\mathbb{R}^{n-1}$ such that for all $w\in\mathbb{S}^{n-2}$,
$\int_{0}^{\pi}\sin\theta\,q(\theta,w)\,d\theta=t\qquad\text{and}\qquad\int_{0}^{\pi}\cos\theta\,q(\theta,w)\,d\theta=\langle x,w\rangle.$
and for all $(\theta,w)\in(0,\pi)\times\mathbb{S}^{n-2}$ and unit tangent vectors $\xi\in T_{w}\mathbb{S}^{n-2}$,
$\begin{split}&\sin^{2}\!\theta\,q(\theta,w)\int_{0}^{\theta}[\sin(\theta-\sigma)\xi^{\top}\nabla^{2}_{\mathbb{S}^{n-2}}q(\sigma,w)\,\xi+\sin\theta\cos\sigma q(\sigma,w)]\,d\sigma\\ &\qquad>\bigg(\int_{0}^{\theta}\sin\sigma\,\xi^{\top}\nabla^{\mathbb{S}^{n-2}}q(\sigma,w)\,d\sigma\bigg)^{\!2}\!.\end{split}$
In this case, the body $K$ is unique up to translations, and $h_{K-v}=h$ for some $v\in\mathbb{R}^{n}$, where
$h(\theta,w)=(n-1)\int_{0}^{\theta}\sin(\theta-\sigma)q(\sigma,w)\,d\sigma,\qquad(\theta,w)\in[0,\pi]\times\mathbb{S}^{n-2}.$
Plan of the article
˜A is proved in Section˜2 and ˜B is proved in Section˜3. The necessary background is given wherever it is needed first.
2. A convolution representation
Our proof of ˜A is inspired by the the Kubota-type formula of Hug, Mussnig, and Ulivelli [Hug2024] in the instance where $i=1$: For a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ and a bounded Borel function $f:\mathbb{S}^{n-1}\to\mathbb{R}$,
@@ 2. A convolution representation
This integral geometric formula expresses the mixed area measure $S_{1}(K,\mathbb{D},{}\cdot{})$ as an average of surface area measures of two-dimensional projections of $K$. The core idea of our argument is to compare ˜2.1 to the disintegration ˜1.5 of the measure $\mu$ to relate the mixed Christoffel problem with the disk to the classical Christoffel problem in two dimensions.
By Theorem˜1.1, we explicitly solve the $2$-dimensional Christoffel problem for the measures $\mu_{E}$, giving bodies $K_{E}\in\mathcal{K}(E)$, $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. This is where the integral kernel
$g_{2}(t)=\sqrt{1-t^{2}}(\pi-\arccos t)+c_{n}t,$
originates, with $c_{n}\in\mathbb{R}$ chosen in such a way to make the function $u\mapsto g_{2}(\left\langle e_{n},u\right\rangle)$ centered on $\mathbb{S}^{n-1}$. If there is a body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that every $K_{E}$ arises as the orthogonal projection $K|E$ of $K$ onto $E$, then we can conclude $\mu=S_{1}(K,\mathbb{D};{{}\cdot{}})$. However, as the subspaces $E$ intersect only in the line spanned by $e_{n}$, compatibility of the $K_{E}$ is no issue; convexity is asserted by condition ˜(iii) in ˜A.
@@
In the following, we first recall some classical facts from convex geometry, in particular on mixed volumes and mixed area measures that will be needed later on. Then we show the existence of the disintegration ˜1.5. In the last section we finally prove ˜A.
@@ 2.1. Preliminaries
First, the Steiner point (see, e.g., [Schneider2014]*p. 50) of a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ is defined as
(2.2) $s(K)=\frac{1}{\kappa_{n}}\int_{\mathbb{S}^{n-1}}h_{K}(u)u\,du\in\mathbb{R}^{n}.$
Next, the mixed volume $V(K_{1},\dots,K_{n})$, $K_{1},\dots,K_{n}\in\mathcal{K}(\mathbb{R}^{n})$ is defined as the suitable coefficient of the homogeneous polynomial obtained by Minkowski addition
@@
with $F(K,u)=\{x\in\mathbb{R}^{n}:h_{K}(u)=\left\langle x,u\right\rangle\}$ the face of $K$ in direction $u\in\mathbb{S}^{n-1}$.
@@
In order to deal with the uniqueness of the solution $K$ to the Christoffel problem involving the disk, we recall Minkowski’s quadratic inequality, a classical result from the Brunn–Minkowski theory. It states that for all convex bodies $K,L,C\in\mathcal{K}(\mathbb{R}^{n})$,
@@
Let $K,L,C\in\mathcal{K}(\mathbb{R}^{n})$ be such that $C-C\subseteq e_{n}^{\perp}$ and $V(K,L,C^{[n-2]})>0$. Then equality holds in ˜2.6 if and only if $\tilde{L}:=\frac{V(K,L,C^{[n-2]})}{V(L,L,C^{[n-2]})}L$ satisfies that $K+F(\tilde{L},e_{n})$ and $\tilde{L}+F(K,e_{n})$ have the same supporting hyperplanes in all $1$-extremal normal directions of $C$.
Observe now that at every point $x\in\mathbb{D}$, the corresponding normal cone $N(\mathbb{D},x)$ is either one- or two-dimensional. This implies that every direction $u\in\mathbb{S}^{n-1}$ is a $1$-extremal normal direction of $\mathbb{D}$, so Theorem˜2.2 specializes for $C=\mathbb{D}$ to the following statement.